Tickets for a concert were sold to adults for $3 and children for $2. If the total receipts were $824 and twice as many adult tickets were sold as childrens tickets how many of each were sold?

Answer by Earlsdon(6294)

Let A = the number of adult tickets sold and C = the number of children's tickets sold.
From the problem description, you are told:
A = 2C "...twice as many adult tickets were sold as children's tickets..."
The cost of the adult tickets can be expressed as ($3)A while the cost of the children's tickets can be expressed as ($2)C, so we can set up the equation for solving as follows:
($3)A+($2)C = $824 You can dispense with the $ signs and just work with the numbers:
3A+2C = 824 Substitute, from above, A = 2C
3A+A = 824
4A = 824
A = 206 and C = A/2 = 103
So 206 adult and 103 children's tickets were sold.